# Overcoming Data Scarcity in Generative Language Modelling for Low-Resource Languages: A Systematic Review

### v1

@@ 3.3.5. Monolingual Models
Papers tagged with the monolingual model category were almost as common as multilingual models in our systematic review, with the approach being adopted by 11% of the papers in this review. These papers encapsulate all studies that trained a generative language model for a particular language. This category differs from monolingual text data as some approaches involve building models for a single LRL without enlisting any form of data augmentation. Some of these approaches include training a Turkish LLM via continual training on various models that were not explicitly trained on Turkish data, such as Mistral-7B (Acikgoz et al., 2024); developing a Tibetan QA corpus generation model that generates text and applies BERT for grammar correction (Sun et al., 2019); building a sequence-to-sequence model for Slovenian as an LRL (Ulčar and Robnik-Šikonja, 2023); creating an Indonesian religious QA model (Rizqullah et al., 2023); and training a Bangla-based generative pre-trained transformer (Salim et al., 2023).
@@ 3.3.6. Adaptive Learning
Adaptive learning is featured as a common method for overcoming data scarcity, appearing in 11% of the papers examined for this systematic review. This method relates to taking pre-existing language models and resuming the training phase of the model to expose it to more data. LLamaTurk is based on the open source LLaMa-7b base and further trained on Turkish text data (Toraman, 2024). A similar approach is done for Turkish using Mistral-7b and GPT-2 as a base model (Acikgoz et al., 2024). Pre-trained models have also been fine-tuned to exploit cross-lingual capabilities for the Basque language (Otegi et al., 2020).
@@ 3.3.9. Miscellaneous
A variety of papers have unique approaches within this review set that have been categorised on their own for tackling the issue of data scarcity when building generative language models. One of these approaches pertains to extending the vocabulary of a large language model by merging the vocabulary of an existing model with the vocabulary of an LRL such as Turkish (Toraman, 2024). Another study adopts the concept of code-mixing, whereby a sequence-to-sequence model is trained on parallel data to create sentences comprising a mix of languages (Bhowmick et al., 2023).
@@ 3.5. RQ4: Which Sources Are Publishing in the Area of Low-Resource Language Modelling?
There are a variety of sources publishing in the area of LRL modelling. Figure 9 indicates that the Association for Computational Linguistics (ACL) is the most active publisher in this area with 35.2% of the papers belonging to their anthology (Sorokin et al., 2022; Agarwal et al., 2022; Gerz et al., 2018; Dandapat and Federmann, 2018; Cahyawijaya et al., 2021; Mao and Yu, 2024; Gao et al., 2020; Marie and Fujita, 2020; Usui and Komiya, 2023; Wang et al., 2018; Gao et al., 2019; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Liu, 2022; Guo et al., 2024b; Wang et al., 2022; Otegi et al., 2020; Scalvini and Debess, 2024). ACL represents 19 different publications spanning 13 conference proceedings, three workshops, and three journals within this review. The Conference on Empirical Methods in Natural Language Processing (Zhang et al., 2020; Cahyawijaya et al., 2021; Wang et al., 2018) and the Transaction of the Association for Computational Linguistics (Marie and Fujita, 2020; Gerz et al., 2018; Wang et al., 2022) are the most frequently cited sources by ACL, both consisting of three papers in this review. Furthermore, the International Conference on Language Resources and Evaluation (Scalvini and Debess, 2024; Otegi et al., 2020) and the Annual Meeting of the Association for Computational Linguistics (Gao et al., 2019, 2020) are other conference papers by ACL that feature twice in this review. ACL also produces individual papers from nine other sources such as the Workshop on Multilingual Information Access (Agarwal et al., 2022); the Workshop on Technologies for Machine Translation of Low-resource Languages (Mao and Yu, 2024); the Conference on Machine Translation (Berckmann and Hiziroglu, 2020); the North American Chapter of the Association for Computational Linguistics (Sorokin et al., 2022); the Workshop on NLP for Similar Languages (Liu, 2022); the International Conference on Computational Linguistics (Guo et al., 2024b); the International Conference on Natural Language Processing for Digital Humanities (Usui and Komiya, 2023); the Annual Conference of the European Association for Machine Translation (Dandapat and Federmann, 2018); the European Chapter of the ACL (Li et al., 2024b).
@@
The Association for Computing Machinery (ACM) is the joint-second most frequently cited publisher in this field, accounting for 18.5% of the papers included in this review (Bhowmick et al., 2023; Yirmibeşoğlu and Güngör, 2023; Maimaiti et al., 2021; Tanwar and Majumder, 2020; Pham Van and Le Thanh, 2022; Wu and Wang, 2023; Mi et al., 2024; Shi et al., 2022; Zhang et al., 2024b; Yu et al., 2020). Although ACM is the joint-second most frequently cited publisher among the sources from this review, its journal publication Transactions on Asian and Low-resource Language Information Processing is the largest individual source of papers in this review, as seen in Figure 10 (Bhowmick et al., 2023; Yirmibeşoğlu and Güngör, 2023; Maimaiti et al., 2021; Tanwar and Majumder, 2020; Mi et al., 2024; Shi et al., 2022; Zhang et al., 2024b; Yu et al., 2020). ACM is also responsible for a single paper via the Symposium on Information and Communication Technology (Pham Van and Le Thanh, 2022) and a single paper with the Asia Pacific Information Technology Conference (Wu and Wang, 2023).
@@ 3.6. RQ5: What Architectures Are Being Used to Produce Low-Resource Language Models?
A diverse set of architectures have been enlisted for LRL generation tasks. For this research question, we have assigned categories to each paper in this review depending on both the general architecture of the models that are built and the specific models that are used. Figure 13 indicates that the transformer architecture dominates as the architecture of choice for 76% of the papers in this review. In terms of the papers that model LRLs, the majority of them employ a transformer-based approach. Recurrent Neural Network-based implementations are the second most frequently adopted architecture for LRL modelling, with nine studies in this systematic review opting to utilise them. The Long Short-Term Memory (LSTM) categorisation applies to six papers. Given that LSTM-based approaches are inherently RNNs, it is worth highlighting that the categorisations are not mutually exclusive. 4% of papers in this review enlist Generative Adversarial Networks (GAN) for building LRL generation models. Additionally, a Gated Recurrent Unit (GRU) approach is implemented by one study, while another study employs a Statistical Machine Translation (SMT) architecture. In terms of the specific models being used to produce LRL models, Figure 14 showcases the most frequently implemented models. GPT-2 appears most often in the studies with 11% of papers enlisting it for LRL modelling. XLM, MT5 and mBERT are each modelled by 7% of studies. T5 is used by 6% papers for LRL modelling. Furthermore, mBART and Mistral-7b are both enlisted by 4% of papers. The rest of the models are used once and include variations of LLama, BERT and BART models.
@@
Text data augmentations have been used by a variety of studies in this review to improve the performance of language generation tasks for LRLs. Back-translation has been found to improve BLEU scores for translation tasks in comparison with no augmentation (Pham et al., 2021; Hong et al., 2024; Berckmann and Hiziroglu, 2020; Yirmibeşoğlu and Güngör, 2023; Nissanka et al., 2020; Dandapat and Federmann, 2018; Pham Van and Le Thanh, 2022). This contrasts slightly with a minor BLEU improvement found by another study that enlists back-translation to improve Cherokee-English machine translation (Zhang et al., 2020). However, the study attributes the limited performance improvement to the lack of in-domain training data. Using a combination of synthetic code-mixed and mixed-script with original data has been found to improve BLEU, METEOR and ROUGE scores for Indic language translation (Bhowmick et al., 2023). Synthesising new text data via paraphrasing as a data augmentation technique has also been shown to increase the BLEU performance of translating from an LRL to English (Guo et al., 2021a). Paraphrasing has improved BLEU results at both a character and word level for question-answering corpus generation for Tibetan (Sun et al., 2019). Paraphrasing has also been used to improve BLEU scores by boosting data in low-resource tasks such as dialogue generation (Gao et al., 2020). Reordering monolingual sentences as a form of text data augmentation has also yielded BLEU score improvements across a baseline of other data augmentation techniques (Wu and Wang, 2023). Similarly, synthesising user-generated text has been used as an augmentation technique to improve BLEU scores for NMT (Marie and Fujita, 2020). Training a language model on a family of related languages is an effective data augmentation technique whereby a study has built an Indonesian language model that achieves similar results to state-of-the-art models with only  20% of the number of parameters (Cahyawijaya et al., 2021). Another approach likens multilingual language modelling of related languages to a 33% increase in the size of the monolingual dataset (Chang et al., 2023). Mass translation as a text data augmentation technique has also been found to improve COPA results for Bengali (Shafayat et al., 2024). Word-level augmentations have been successful at improving BLEU scores across other augmentation techniques for low-resource NMT (Maimaiti et al., 2021; Wang et al., 2018; Gao et al., 2019).
@@ 3.8. RQ7: How Scarce Is the Data in Studies Tackling Low-Resource Language Modelling?
As part of this review, we determine the quantity of data that has been used to model generative language models for LRLs. This process involved investigating and exploring the corpora that have been used to train models for LRL in this review. Given the ambiguity surrounding corpus statistics in many of the papers reviewed, this research question focuses on whether the study specifies the number of sentences used to train the LRL model. We found that three different datasets were used to train generative language models for Khmer (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022) with data ranging from 70k-579k sentences. Figure 15 highlights the languages with a single data source, where no additional sources are available. This visualisation indicates that Minnan has been modelled with the largest sentence count of 1.2m sentences (Gerz et al., 2018). From this aggregation, a count of 390k sentences is the average count for training language models for these LRLs. In several studies, data is quantified in terms of size in gigabytes rather than sentence count. For instance, Bengali models were trained on 26GB of data (Salim et al., 2023), while Estonian, Turkish, and Indonesian models used 6.1GB (Downey et al., 2024), 180GB (Acikgoz et al., 2024), and 23.8GB (Cahyawijaya et al., 2021) of data, respectively.
@@ 3.9. RQ8: How Are the Models in These Studies Being Evaluated?
Numerous evaluation methods have been enlisted to assess the performance of NLG models for LRLs in this review. Figure 16 indicates that BLEU dominates as the evaluation method of choice with 61% of papers using it to gauge the efficacy of their models (Salim et al., 2023; Bhowmick et al., 2023; Tanwar and Majumder, 2020; Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Cahyawijaya et al., 2021; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Wu and Wang, 2023; Sun et al., 2019; Guo et al., 2021a; Gao et al., 2020; Marie and Fujita, 2020; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Yu et al., 2020; Phan-Vu et al., 2017; Li et al., 2022a; Otegi et al., 2020; Zhang et al., 2024a). ROUGE is used by 9% of studies (Bhowmick et al., 2023; Saini and Vidhyarthi, 2023; Holiel et al., 2023; Li et al., 2022a), hence it is the second most frequently used evaluation method in this review. Perplexity, chrF, and METEOR are each used in 7% of the papers to evaluate their models (Salim et al., 2023; Chang et al., 2023; Bhowmick et al., 2023; Maimaiti et al., 2021; Toraman, 2024; Saini and Vidhyarthi, 2023; Holiel et al., 2023; Li et al., 2022a; Scalvini and Debess, 2024; Zhang et al., 2024a; Vo and Tran, 2023). Furthermore, five evaluation methods each appear in 6% of the studies. These methods include human evaluation (Niyogi and Bhattacharya, 2024; Dandapat and Federmann, 2018; Hong et al., 2024); chrF++ (Mao and Yu, 2024; Li et al., 2024b; Guo et al., 2024b), sacreBLEU (Wongso et al., 2023; Hong et al., 2024; Liu, 2022); COMET (Mao and Yu, 2024; Hong et al., 2024; Guo et al., 2024b); and F1-score (Sorokin et al., 2022; Rizqullah et al., 2023; Otegi et al., 2020).
@@ 3.10. RQ9: What Is the NLG Task?
A variety of generative models have been created for LRLs by studies in this review. Figure 17 showcases each NLG task by frequency. Given that some papers attempt to build models for multiple generative tasks, the categorisations are not mutually exclusive. Translation is the most popular generative task category in this review, with 57% of the studies focusing on the task (Bhowmick et al., 2023; Agarwal et al., 2022; Gerz et al., 2018; Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Toraman, 2024; Acikgoz et al., 2024; Yirmibeşoğlu and Güngör, 2023; Cahyawijaya et al., 2021; Wongso et al., 2023; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Vo and Tran, 2023; Downey et al., 2024; Pham et al., 2021; Haque et al., 2021; Shi et al., 2022; Wu and Wang, 2023; Sun et al., 2019; Guo et al., 2021a; Wang et al., 2018; Li et al., 2024b; Yu et al., 2020; Phan-Vu et al., 2017; Guo et al., 2024b; Wang et al., 2022; Otegi et al., 2020; Zhang et al., 2024a). Furthermore, general language modelling is the second most popular language generation task in this review with 24% of the papers tackling it (Niyogi and Bhattacharya, 2024; Salim et al., 2023; Chang et al., 2023; Gerz et al., 2018; Toraman, 2024; Acikgoz et al., 2024; Cahyawijaya et al., 2021; Downey et al., 2024; Berckmann and Hiziroglu, 2020; Holiel et al., 2023; Li et al., 2022a; Ulčar and Robnik-Šikonja, 2023). We categorise general language modelling as research relating to building large pretrained models as a base for various downstream tasks. Question-answering is also somewhat popular with 15% of the papers building models for the task (Shafayat et al., 2024; Sorokin et al., 2022; Agarwal et al., 2022; Sun et al., 2019; Rizqullah et al., 2023; Otegi et al., 2020; Vo and Tran, 2023). Summarisation (Gao et al., 2020) and dialogue generation models are also modelled by a single paper in this systematic review (Holiel et al., 2023).

### v2

@@ 3.3.5. Monolingual Models
Papers tagged with the monolingual model category were almost as common as multilingual models in our systematic review, with the approach being adopted by 11% of the papers. These papers encapsulate all studies that trained a generative language model for a particular language. This category differs from monolingual text data as some approaches involve building models for a single LRL without using any form of data augmentation. Some of these approaches include training a Turkish LLM via continual training on various models that were not explicitly trained on Turkish data, such as Mistral-7B (Acikgoz et al., 2024); developing a Tibetan question-answering (QA) corpus generation model that generates text and applies BERT for grammar correction (Sun et al., 2019); building a sequence-to-sequence model for Slovenian as an LRL (Ulčar and Robnik-Šikonja, 2023); creating an Indonesian religious QA model (Rizqullah et al., 2023); and training a Bangla-based generative pre-trained transformer (Salim et al., 2023).
@@ 3.3.6. Adaptive Learning
Adaptive learning is featured as a common method for overcoming data scarcity, appearing in 11% of the papers examined for this systematic review. This method relates to taking pre-existing language models and resuming the training phase of the model to expose it to more data. LLamaTurk is based on the open-source LLaMa-7b base and further trained on Turkish text data (Toraman, 2024). A similar approach is done for Turkish using Mistral-7b and GPT-2 as a base model (Acikgoz et al., 2024). Pre-trained models have also been fine-tuned to exploit cross-lingual capabilities for the Basque language (Otegi et al., 2020).
@@ 3.3.9. Miscellaneous
A variety of studies have unique approaches that have been categorised on their own for tackling the issue of data scarcity when building generative language models. One of these approaches pertains to extending the vocabulary of a large language model by merging the vocabulary of an existing model with the vocabulary of an LRL such as Turkish (Toraman, 2024). Another study adopts the concept of code-mixing, whereby a sequence-to-sequence model is trained on parallel data to create sentences comprising a mix of languages (Bhowmick et al., 2023).
@@ 3.5. RQ4: Which Sources Are Publishing in the Area of Low-Resource Language Modelling?
There are a variety of sources publishing in the area of LRL modelling. Figure 9 indicates that the Association for Computational Linguistics (ACL) is the most active publisher in this area with 35.2% of the papers belonging to their anthology (Sorokin et al., 2022; Agarwal et al., 2022; Gerz et al., 2018; Dandapat and Federmann, 2018; Cahyawijaya et al., 2021; Mao and Yu, 2024; Gao et al., 2020; Marie and Fujita, 2020; Usui and Komiya, 2023; Wang et al., 2018; Gao et al., 2019; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Liu, 2022; Guo et al., 2024b; Wang et al., 2022; Otegi et al., 2020; Scalvini and Debess, 2024). ACL represents 19 different publications spanning 13 conference proceedings, three workshops, and three journals. The Conference on Empirical Methods in Natural Language Processing (Zhang et al., 2020; Cahyawijaya et al., 2021; Wang et al., 2018) and the Transaction of the Association for Computational Linguistics (Marie and Fujita, 2020; Gerz et al., 2018; Wang et al., 2022) are the most frequently cited sources by ACL, both consisting of three papers. Furthermore, the International Conference on Language Resources and Evaluation (Scalvini and Debess, 2024; Otegi et al., 2020) and the Annual Meeting of the Association for Computational Linguistics (Gao et al., 2019, 2020) are other conference papers by ACL that feature twice in this review. ACL also produces individual papers from nine other sources such as the Workshop on Multilingual Information Access (Agarwal et al., 2022); the Workshop on Technologies for Machine Translation of Low-resource Languages (Mao and Yu, 2024); the Conference on Machine Translation (Berckmann and Hiziroglu, 2020); the North American Chapter of the Association for Computational Linguistics (Sorokin et al., 2022); the Workshop on NLP for Similar Languages (Liu, 2022); the International Conference on Computational Linguistics (Guo et al., 2024b); the International Conference on Natural Language Processing for Digital Humanities (Usui and Komiya, 2023); the Annual Conference of the European Association for Machine Translation (Dandapat and Federmann, 2018); the European Chapter of the ACL (Li et al., 2024b).
@@
The Association for Computing Machinery (ACM) is the joint-second most frequently cited publisher in this field, accounting for 18.5% of the papers (Bhowmick et al., 2023; Yirmibeşoğlu and Güngör, 2023; Maimaiti et al., 2021; Tanwar and Majumder, 2020; Pham Van and Le Thanh, 2022; Wu and Wang, 2023; Mi et al., 2024; Shi et al., 2022; Zhang et al., 2024b; Yu et al., 2020). Although ACM is the joint-second most frequently cited publisher among the sources from this review, its journal publication Transactions on Asian and Low-resource Language Information Processing is the largest individual source of papers, as seen in Figure 10 (Bhowmick et al., 2023; Yirmibeşoğlu and Güngör, 2023; Maimaiti et al., 2021; Tanwar and Majumder, 2020; Mi et al., 2024; Shi et al., 2022; Zhang et al., 2024b; Yu et al., 2020). ACM is also responsible for a single paper via the Symposium on Information and Communication Technology (Pham Van and Le Thanh, 2022) and a single paper with the Asia Pacific Information Technology Conference (Wu and Wang, 2023).
@@ 3.6. RQ5: What Architectures Are Being Used to Produce Low-Resource Language Models?
A diverse set of architectures have been adopted for LRL generation tasks. For this research question, we have assigned categories to each paper depending on both the general architecture of the models that are built and the specific models that are used. Figure 13 indicates that the transformer architecture dominates as the architecture of choice for 76% of the papers. In terms of the papers that model LRLs, the majority of them employ a transformer-based approach. Recurrent Neural Network-based implementations are the second most frequently adopted architecture for LRL modelling, with nine studies in this systematic review opting to utilise them. The Long Short-Term Memory (LSTM) categorisation applies to six papers. Given that LSTM-based approaches are inherently RNNs, it is worth highlighting that the categorisations are not mutually exclusive. 4% of papers use Generative Adversarial Networks (GAN) for building LRL generation models. Additionally, a Gated Recurrent Unit (GRU) approach is implemented by one study, while another study employs a Statistical Machine Translation (SMT) architecture. In terms of the specific models being used to produce LRL models, Figure 14 showcases the most frequently implemented models. GPT-2 appears most often in the studies with 11% of papers adapting it for LRL modelling. XLM, MT5 and mBERT are each modelled by 7% of studies. T5 is used by 6% papers for LRL modelling. Furthermore, mBART and Mistral-7b are both used by 4% of papers. The rest of the models are used once and include variations of LLama, BERT and BART models.
@@
Text data augmentations have been used by a variety of studies to improve the performance of language generation tasks for LRLs. Back-translation has been found to improve BLEU scores for translation tasks in comparison with no augmentation (Pham et al., 2021; Hong et al., 2024; Berckmann and Hiziroglu, 2020; Yirmibeşoğlu and Güngör, 2023; Nissanka et al., 2020; Dandapat and Federmann, 2018; Pham Van and Le Thanh, 2022). This contrasts slightly with a minor BLEU improvement found by another study that uses back-translation to improve Cherokee-English machine translation (Zhang et al., 2020). However, the study attributes the limited performance improvement to the lack of in-domain training data. Using a combination of synthetic code mixed and mixed-script with original data has been found to improve BLEU, METEOR and ROUGE scores for Indic language translation (Bhowmick et al., 2023). Synthesising new text data via paraphrasing as a data augmentation technique has also been shown to increase the BLEU performance of translating from an LRL to English (Guo et al., 2021a). Paraphrasing has improved BLEU results at both a character and word level for QA corpus generation for Tibetan (Sun et al., 2019). Paraphrasing has also been used to improve BLEU scores by boosting data in low-resource tasks such as dialogue generation (Gao et al., 2020). Reordering monolingual sentences as a form of text data augmentation has also yielded BLEU score improvements across a baseline of other data augmentation techniques (Wu and Wang, 2023). Similarly, synthesising user-generated text has been used as an augmentation technique to improve BLEU scores for NMT (Marie and Fujita, 2020). Training a language model on a family of related languages is an effective data augmentation technique whereby a study has built an Indonesian language model that achieves similar results to state-of-the-art models with only  20% of the number of parameters (Cahyawijaya et al., 2021). Another approach likens multilingual language modelling of related languages to a 33% increase in the size of the monolingual dataset (Chang et al., 2023). Mass translation as a text data augmentation technique has also been found to improve COPA results for Bengali (Shafayat et al., 2024). Word-level augmentations have been successful at improving BLEU scores across other augmentation techniques for low-resource NMT (Maimaiti et al., 2021; Wang et al., 2018; Gao et al., 2019).
@@ 3.8. RQ7: How Scarce Is the Data in Studies Tackling Low-Resource Language Modelling?
As part of this review, we determine the quantity of data that has been used to model generative language models for LRLs. This process involved investigating and exploring the corpora that have been used to train models for LRLs. Given the ambiguity surrounding corpus statistics in many of the papers reviewed, this research question focuses on whether the study specifies the number of sentences used to train the LRL model. We found that three different datasets were used to train generative language models for Khmer (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022) with data ranging from 70k-579k sentences. Figure 15 highlights the languages with a single data source, where no additional sources are available. This visualisation indicates that Minnan has been modelled with the largest sentence count of 1.2m sentences (Gerz et al., 2018). From this aggregation, a count of 390k sentences is the average count for training language models for these LRLs. In several studies, data is quantified in terms of size in gigabytes rather than sentence count. For instance, Bengali models were trained on 26GB of data (Salim et al., 2023), while Estonian, Turkish, and Indonesian models used 6.1GB (Downey et al., 2024), 180GB (Acikgoz et al., 2024), and 23.8GB (Cahyawijaya et al., 2021) of data, respectively.
@@ 3.9. RQ8: How Are the Models in These Studies Being Evaluated?
Numerous evaluation methods have been adopted to assess the performance of NLG models for LRLs. Figure 16 indicates that BLEU dominates as the evaluation method of choice with 61% of papers using it to gauge the efficacy of their models (Salim et al., 2023; Bhowmick et al., 2023; Tanwar and Majumder, 2020; Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Cahyawijaya et al., 2021; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Wu and Wang, 2023; Sun et al., 2019; Guo et al., 2021a; Gao et al., 2020; Marie and Fujita, 2020; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Yu et al., 2020; Phan-Vu et al., 2017; Li et al., 2022a; Otegi et al., 2020; Zhang et al., 2024a). ROUGE is used by 9% of studies (Bhowmick et al., 2023; Saini and Vidhyarthi, 2023; Holiel et al., 2023; Li et al., 2022a), hence it is the second most frequently used evaluation method. Perplexity, chrF, and METEOR are each used in 7% of the papers to evaluate their models (Salim et al., 2023; Chang et al., 2023; Bhowmick et al., 2023; Maimaiti et al., 2021; Toraman, 2024; Saini and Vidhyarthi, 2023; Holiel et al., 2023; Li et al., 2022a; Scalvini and Debess, 2024; Zhang et al., 2024a; Vo and Tran, 2023). Furthermore, five evaluation methods each appear in 6% of the studies. These methods include human evaluation (Niyogi and Bhattacharya, 2024; Dandapat and Federmann, 2018; Hong et al., 2024); chrF++ (Mao and Yu, 2024; Li et al., 2024b; Guo et al., 2024b), sacreBLEU (Wongso et al., 2023; Hong et al., 2024; Liu, 2022); COMET (Mao and Yu, 2024; Hong et al., 2024; Guo et al., 2024b); and F1-score (Sorokin et al., 2022; Rizqullah et al., 2023; Otegi et al., 2020).
@@ 3.10. RQ9: What Is the NLG Task?
A variety of generative models have been created for LRLs by studies. Figure 17 showcases each NLG task by frequency. Given that some papers attempt to build models for multiple generative tasks, the categorisations are not mutually exclusive. Translation is the most popular generative task category, with 57% of the studies focusing on the task (Bhowmick et al., 2023; Agarwal et al., 2022; Gerz et al., 2018; Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Toraman, 2024; Acikgoz et al., 2024; Yirmibeşoğlu and Güngör, 2023; Cahyawijaya et al., 2021; Wongso et al., 2023; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Vo and Tran, 2023; Downey et al., 2024; Pham et al., 2021; Haque et al., 2021; Shi et al., 2022; Wu and Wang, 2023; Sun et al., 2019; Guo et al., 2021a; Wang et al., 2018; Li et al., 2024b; Yu et al., 2020; Phan-Vu et al., 2017; Guo et al., 2024b; Wang et al., 2022; Otegi et al., 2020; Zhang et al., 2024a). Furthermore, general language modelling is the second most popular language generation task with 24% of the papers tackling it (Niyogi and Bhattacharya, 2024; Salim et al., 2023; Chang et al., 2023; Gerz et al., 2018; Toraman, 2024; Acikgoz et al., 2024; Cahyawijaya et al., 2021; Downey et al., 2024; Berckmann and Hiziroglu, 2020; Holiel et al., 2023; Li et al., 2022a; Ulčar and Robnik-Šikonja, 2023). We categorise general language modelling as research relating to building large pre-trained models as a base for various downstream tasks. Question-answering is also somewhat popular with 15% of the papers building models for the task (Shafayat et al., 2024; Sorokin et al., 2022; Agarwal et al., 2022; Sun et al., 2019; Rizqullah et al., 2023; Otegi et al., 2020; Vo and Tran, 2023). Summarisation (Gao et al., 2020) and dialogue generation models are also modelled by a single paper in this systematic review (Holiel et al., 2023).